OpenStudy (anonymous):

Question Part Points Submissions Used A model for the population in a small community after t years is given by the equation P(t) = P0ekt, k > 0. (a) If the initial population has doubled in 5 years, how long will it take to triple? (Round your answer to two decimal places.) yr If the initial population has doubled in 5 years, how long will it take to quadruple? (Round your answer to the nearest integer.) yr (b) If the population of the community in part (a) is 10,000 after 3 years, what was the initial population? (Round your answer to the nearest integer.) P0 =

OpenStudy (anonymous):

Part A: \[\frac{P(5)}{P_o} = 2 = e^{5k} \rightarrow \ln(2) = 5k \rightarrow k = \frac{\ln(2)}{5} \approx 0.1386\] \[P(t) = P_oe^{0.1386t}\] so to find the time to triple, \[\frac{P(t)}{P_o} = 3 = e^{0.1386t} \rightarrow \ln(3) = 0.1386t \rightarrow t = 7.92 years\] \[\frac{P(t)}{P_o} = 4 = e^{0.1386t} \rightarrow \ln(4) = 0.1386t \rightarrow t = 10 years\] Part B: \[P_o = P(t)e^{-0.1386t} = 10,000e^{-0.1386*3} = 6,598\]

OpenStudy (anonymous):

THANKS!